OpenStudy (samsan9):

If your pour a cup of coffee that is 200^0 Fahrenheit and set it on a desk in a room that iis 68 Fahrenheit, and 10 minutes later it is 145 Fahrenheit, what temperature will it be 15 minutes after you originally poured it

OpenStudy (anonymous):

not as bad as it looks ready?

OpenStudy (samsan9):

yes

OpenStudy (anonymous):

always work with the temperature differences \[200-68=132\] and \[145-68=77\]

OpenStudy (samsan9):

ok

OpenStudy (anonymous):

then without being fancy, you can model the temperature difference as \[\large 132\times \left(\frac{77}{132}\right)^{\frac{t}{10}}\]

OpenStudy (anonymous):

notice i used only the numbers given to you, or rather computed the differences with those numbers i did not mess around with \[T=T_0+e^{-kt}\] or whatever

OpenStudy (samsan9):

yeah i don''t really like that formula especially since i always forget how to find t

OpenStudy (anonymous):

since you want the temperature in fifteen minutes, replace \(t\) by \(15\) and compute \[132\times \left(\frac{77}{132}\right)^{\frac{15}{10}}\] or \[132\times \left(\frac{77}{132}\right)^{\frac{3}{2}}\]

OpenStudy (anonymous):

then don't forget to add \(68\) to your answer, because that is the difference in the temperatures and you have to add back on the room temp

OpenStudy (anonymous):

looks like \(126.8\) http://www.wolframalpha.com/input/?i=132%2877%2F132%29^%283%2F2%29%2B68